OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16pi sq. in a)32/3)pi cubic inches b)8pi cubic inches c)4/3)pi cubic inches

OpenStudy (sparrow2):

16p=4pr^2 so r=2 v=4/3pr^3 so v=4/3xpx8=32/3p cubic inch so a,i think i answered it before :D

OpenStudy (anonymous):

lol Thank you sparrow!